KQuestion 3 (of 12) 3. value: 10.00 points 3 attempts left Check my work Enter your answer in the provided box. At 46°C a sample of ammonia gas exerts a pressure of 3.01 atm. What is the pressure when the volume of the gas is reduced to one-fifth of the original value at the same temperature? atm eBook & Resources References
media%2F391%2F39134db9-eebe-4bf6-b7d6-53
Answer #1

1) From Boyle's Law:
P1V1 = P2V2
Given that, P1 = 3.01 atm
Assuming Initial volume = 1 L
New volume = 1/5

(3.01 atm)(1) = (P2)(1/5)
P2 = 3.01*5 atm
P2 = 15.05 atm

2) From Charles's Law:
V1/T1 = V2/T2
Given that, T1 = 99 oC
Initial volume =9.30 L
Final volume = 1.90 L

9.30/99 = 1.90/T2
T2 = 20.22 oC